OpenStudy (anonymous):

What is the constant of variation? An x−y table shows the following coordinate pairs: two comma twenty-four, four comma twelve, and six comma eight.

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (danjs):

The constant of variation is the slope of the line containing those 2 points. Y=k*x, where k is the slope and the constant of variation

OpenStudy (anonymous):

I could use anyone

OpenStudy (danjs):

any 2 points yes

OpenStudy (anonymous):

Are you saying the answer is 2

OpenStudy (danjs):

no, i said, you can use any 2 of those 3 points and calculate the slope

OpenStudy (danjs):

(2,24) , (4, 12) and (6, 8)

OpenStudy (anonymous):

@AnimeAsgardian help I don't get it

OpenStudy (danjs):

Let two of those points be - (x,y) and (x1,y1) , find \[k = \frac{ y - y1 }{ x - x1 }\]

OpenStudy (danjs):

plug in the numbers

OpenStudy (animeasgardian):

He left o-o

OpenStudy (anonymous):

How rude of him!!!!!!!!!

OpenStudy (anonymous):

Can you help?

OpenStudy (animeasgardian):

I'll try

OpenStudy (anonymous):

It is a weird question

OpenStudy (animeasgardian):

Yeah

OpenStudy (anonymous):

WHY is it saying i am typing a reply when i am not doing anything

OpenStudy (animeasgardian):

haha

OpenStudy (anonymous):

So can you help

OpenStudy (anonymous):

Is it 2??

OpenStudy (animeasgardian):

2?

OpenStudy (animeasgardian):

Are there choices?

OpenStudy (anonymous):

OR 12??

OpenStudy (animeasgardian):

What are the answers? Is it multiple choice?

OpenStudy (anonymous):

48 24 12 2

OpenStudy (animeasgardian):

48 24 12 2 ok....

OpenStudy (anonymous):

After these other question i have other things to do which really sucks

OpenStudy (animeasgardian):

I'm not sure what that answer is sorry...

OpenStudy (anonymous):

i am probably going to be up all night :(

OpenStudy (anonymous):

Thats ok

OpenStudy (animeasgardian):

Is there another question I could help you with?

OpenStudy (anonymous):

Yes

OpenStudy (animeasgardian):

Ok give me all of the questions

OpenStudy (anonymous):

hold on not letting me paste it

OpenStudy (animeasgardian):

Okai

OpenStudy (anonymous):

Which of the following is not true? Which of the following is not true? The origin is at the point zero comma zero. The y-coordinate determines the vertical distance from the origin. A coordinate plane is divided into four quadrants. An origin pair shows the horizontal and vertical distance of a point from the origin.

OpenStudy (anonymous):

@AnimeAsgardian

OpenStudy (animeasgardian):

the first one is correct so we can scratch that one out

OpenStudy (danjs):

did you calculate the slope?

OpenStudy (anonymous):

yes @DanJS

OpenStudy (anonymous):

Okay stratch the first one of

OpenStudy (animeasgardian):

stretch the 3rd one too

OpenStudy (anonymous):

Ok

OpenStudy (danjs):

(8-12) / (6-4) = -4/2 = -2 constant of variation is -2

OpenStudy (animeasgardian):

ugh i can't spell sorry XD

OpenStudy (anonymous):

That's ok

OpenStudy (animeasgardian):

hmm...

OpenStudy (anonymous):

@DanJS on what question ??

OpenStudy (danjs):

the original one you asked, constant of variation = -2

OpenStudy (anonymous):

There is no -2

OpenStudy (anonymous):

@AnimeAsgardian more help please

OpenStudy (danjs):

Oh i see,

OpenStudy (danjs):

Constant of variation k in y = kx use point (x,y) = (2,24) 24 = k*2 k = 24/2 = 12

OpenStudy (anonymous):

Do you think it will be faster for me to finish homework if i log out

OpenStudy (anonymous):

Thank you

OpenStudy (animeasgardian):

the one you asked me is the last one

OpenStudy (anonymous):

Thnx

OpenStudy (animeasgardian):

Any more questions?

OpenStudy (anonymous):

Yes

OpenStudy (animeasgardian):

o_o

OpenStudy (anonymous):

@DanJS what

OpenStudy (anonymous):

hello any one home :) lol

OpenStudy (danjs):

That is the question you posted, which one is FALSE,

OpenStudy (anonymous):

I know it's the last one

OpenStudy (danjs):

no it is the second one, The y coordinate determined the distance from the x-axis , not the origin.

OpenStudy (anonymous):

If there is an inverse relationship between x and y, and the constant of variation is forty-two, what is y when x equals three?

OpenStudy (anonymous):

@DanJS there is no The y coordinate determined the distance from the x-axis , not the origin.

OpenStudy (anonymous):

I think 9

OpenStudy (anonymous):

Can you help

OpenStudy (danjs):

inverse relationship means y = k /x

OpenStudy (anonymous):

I know is the last one i asked 9 though

OpenStudy (anonymous):

hello

OpenStudy (danjs):

y = 42 / 3

OpenStudy (anonymous):

14

OpenStudy (danjs):

yes

OpenStudy (anonymous):

Thank you next question??

OpenStudy (danjs):

got to go ,,

OpenStudy (anonymous):

ok

OpenStudy (animeasgardian):

What's the next question?